A square cardboard of side 20m has a 4m wide border round three of its sides. Find the area of the border.
Answer by edjones(4080)

Let's give the border all 4 sides to begin. That makes it 28m square.
A=s^2=28^2=784m^2
For the original square: A=20^2=400
784-400=384 area of border with 4 sides
By removing one of the sides we must subtract 28*4=112
384-112=272m^2 Answer.
.
Ed